OpenStudy (anonymous):

Amanda read 15 pages of her 150-page book in 2 hours. At this rate, how long will it take her to read the entire book? a. 2 hours b. 5 hours c. 10 hours d. 20 hours @dtan5457

OpenStudy (some.random.cool.kid):

well thats easy, just divde

OpenStudy (some.random.cool.kid):

if its 15 per /2 h

OpenStudy (some.random.cool.kid):

so for instance 30 pages will be 4 hrs

OpenStudy (danjs):

pages = k*(#hours)

OpenStudy (some.random.cool.kid):

and 45 will be 6 hrs so theres a great enough example i made it plain and simple

OpenStudy (danjs):

use the point (2 , 150), and find the constant k, for the line, through the origin

OpenStudy (some.random.cool.kid):

also you can multiply 15 pages per hrs in your answer to see what you get

OpenStudy (danjs):

use the point (2 , 15) pages = k*hours 15 = k* 2 k = 15/2 pages per hour pages = (15/2)*(hours)

OpenStudy (danjs):

150 = (15/2) * (hours) hours = ...

OpenStudy (danjs):

\[hours = \frac{ 150 * 2 }{ 15 }\]

OpenStudy (some.random.cool.kid):

well ill just go up to 60 pages... 60 is 8 hrs there just keep that going or do what dan said above as another method which works for you :}

OpenStudy (anonymous):

it think the answer is 20 hours

OpenStudy (some.random.cool.kid):

yep

OpenStudy (danjs):

the rate of reading is 15/2 pages per hour it is a direct variation y = (15/2)x where y=pages and x=hours let y = 150 pages, solve for x, you get 20 , in case you want to know how to go about it